10:00 We are currently experiencing a service outage in London and are investigating. Updates to follow.
10:30 We have opened a remote support case with the hosting provider to investigate a suspected power fault to our equipment.
11:45 We have initialised DR procedures to restore services in an alternative Data Centre. Updates to follow
13:00 The hosting provider investigating and found that both A and B power feeds to the cabinet had tripped. Power has been restored to the cabinet.
We apologies for the inconvenience caused.